7. z-scores and standardized scores Is a z-score a standardized score? No Yes Consider the following distribution of scores with a mean of 50 and a standard deviation of 10. For the letters A, B, C, and D in the boxes beneath the ine labeled z give the z-scores corresponding to each position in the distribution. One z-score is already filled in (-1) Suppose you also want to standardize these scores to a k scale where the mean of k is 100 and the standard deviation is 20·For the letters E, F, G, and H in the boxes beneath the line labeled k give the k-scores corresponding to each position in the distribution. One k-score is already filled in (120)30 40 50 60 70 A 1
Answer #1

A z-score is indeed a standardized score because a standardized score shows the number of standard deviation the data valuesFor B, the value x = 50. Thus, the corresponding Z-score is, x-u 2=- 50 - 50 10 = 0 For C, the value x = 60. Thus, the corresNow the z- scores obtained above are to standardize to a “k” scale with a mean 100 and a standard deviation 20. Thus, u = 100For the value of z-score B=0, the value of x is, ==x-u x=100+20(0) x=100 For the value of z-score C =1, the value of x is, X-Thus, the final graph showing all obtained values is as follows: 30 40 50 60 70 TE C=1 D=2 A=-2 60 -1 80 B=0 100 120 140 . He